# An Overview on Over-the-air Electromagnetic Signal Processing

### v2

@@ V-C1 Reconfigurable Intelligent Surfaces
Specifically, by applying specific phase shifts to the incident EM waves, RISs can reflect signals in directions that deviate from the conventional Snell’s law. This phenomenon, referred to as the universal Snell’s law, allows RISs to impose arbitrary directionality on reflected signals, resulting in highly customizable beamforming and beam focusing. This capability enhances signal strength and reception, and depending on the specific hardware structure and reconfigurable elements of the RIS, additional functionalities such as beam collimation, beam splitting, polarization control, and various forms of analog processing can also be achieved [23, 24, 22, 25]. This versatility positions RISs as a promising solution for addressing the complex challenges faced in modern wireless communication systems. Applications of RISs span communication, sensing, localization/tracking, and the detection of passive objects. Ongoing research is also investigating new avenues for leveraging RISs in mapping and imaging to reconstruct high-fidelity and high-resolution maps of the surrounding radio environment [55].
@@
As an example of the basic functionality of a RIS, let us consider a transmitting antenna, e.g., an access point (AP), that wants to transmit to a user equipment (UE) that is not in direct visibility due to obstruction raised by the presence of an obstacle, e.g., in a configuration similar to Fig. 11. Thanks to the deployment of the RIS, it is possible to create a virtual line-of-sight (LOS) link between the AP and the UE. Specifically, we consider a square RIS consisting of $K$ small unit cells (i.e., scattering elements) spaced $\lambda/2$ apart providing reconfigurable phase shifts, each being characterized by reflection coefficients $\rho_{i,j}=e^{\jmath\theta_{i,j}}$, with $i,j=1,2,\dots,\sqrt{K}$ being the 2D position indexes. According to the discussion in Sec. III-C, a conventional RIS with uncoupled elements is characterized by a diagonal reflection matrix ${\bf{R}}(\bm{\theta})={\rm diag}\left(r_{1},r_{2},\ldots,r_{K}\right)$ (diagonal RIS), where $r_{k}=\rho_{i,j}$ is the reflection coefficient of the $k$-th element at position $(i,j)$, where $k=i+\sqrt{K}(j-1)$, with $i,j=1,2,\dots,\sqrt{K}$.
@@
$\theta_{i,j}=-\pi i\left(u_{x}\left(\Theta_{\mathrm{inc}}\right)+u_{x}(\Theta_{\mathrm{ref}})\right)-\pi j\left(u_{y}\left(\Theta_{\mathrm{inc}}\right)+u_{y}(\Theta_{\mathrm{ref}})\right)$ (44)
@@
One important issue, often overlooked in the literature, arises from reflections in unwanted directions caused by uncontrolled RF sources in the environment. These reflections can generate interference, which may be challenging to mitigate. Diagonal RISs may be ineffective in addressing this problem due to the limited available number of DoF in their reflection matrix, which is equal to $K$. A potential alternative is the use of non-diagonal RISs, which allow controllable coupling between elements. This configuration significantly increases the available number of DoF, reaching up to $K\,(K-1)/2$ in reciprocal structures [45]. An even greater number of DoF (up to $K^{2}$) can be achieved with non-reciprocal designs. Interestingly, a non-diagonal RIS, as a 2D structure, could theoretically achieve the same performance as a SIM, which is a 3D structure but with 2D input and output surfaces. However, as both technologies are still in their early stages, a detailed comparison in terms of implementation complexity is not yet available.
@@ V-C2 Self-Conjugating Metasurfaces
We now present an example of iterative over-the-air EM processing involving two devices located within their respective radiative regions. The objective is to achieve joint communication and optimal beamforming in a MIMO link without requiring channel state information (CSI) estimation or signaling, thereby reducing the latency and overhead typically associated with large MIMO systems. The idea is to exploit an EM device at one end of the link that performs the complex conjugation of the signal coming from the other device, thereby achieving retrodirectivity. Due to this processing capability, such a device is referred to as self-conjugating metasurface (SCM). Additionally, it is envisioned that the SCM introduces appropriate phase shifts, uniformly across all cells, into the retransmitted signals, thereby incorporating information through phase modulation,
This dual-function device, referred to as a modulating SCM in the following, can be considered an evolution of retrodirective antenna arrays, which are systems capable of reflecting incoming signals back toward their direction of arrival [56]. While retrodirective antennas have been studied for decades, the feasibility of retrodirective metasurfaces has only recently been demonstrated in the literature [57]. By leveraging modulating SCMs, it is possible to devise a communication scheme that involves an iterative exchange of signals between a full-duplex central unit equipped with a conventional antenna array with $N$ elements, such as an AP, and a device equipped with a modulating SCM with $M$ elements, such as a sensor, which has queued data intended for the AP. The former continuously transmits an interrogation signal while the latter retro-directs the received signal after modulating it according to the data addressed to the interrogation device.
@@
0: Initialization: generate a guess unit norm beamforming vector ${\bf{x}}[0]$ ;
@@
       2: transmit: $\sqrt{P_{\text{T}}}\,{\bf{x}}[k-1]$ ;
@@
       3: ${\bf{z}}[k]=\sqrt{P_{\text{T}}}\,{\bf{H}}\,{\bf{x}}[k-1]+{\bm{\eta}}[k]$ ;
@@
       4: ${\bf{r}}[k]=ge^{\jmath\theta[k]}{\bf{z}}^{*}[k]$ ;
@@
       5: receive: ${\bf{y}}[k]=e^{\jmath\theta[k]}\,{\bf{A}}^{*}\,{\bf{x}}^{*}[k-1]+{\bf{n}}^{*}[k]$ ;
@@
       6: ${\bf{x}}[k]={{\bf{y}}^{*}[k]}/{\left\|{\bf{y}}[k]\right\|}$ ;
@@
The process begins (step 0 of the pseudo-code) with the AP generating a beamforming vector, ${\bf{x}}[0]\in\mathbb{C}^{1\times N}$. Since the position of the sensor is unknown at startup, this initial beamforming vector is chosen randomly (with norm one) to allow the signal transmitted using ${\bf{x}}[0]$ to cover multiple directions, as shown in Fig. 13, first beam pattern. Obviously, any prior information (e.g., past transmission, the position of the sensor, etc.) can be exploited to speed up the estimation of the optimal beamforming vector. An iterative process is then initiated (see Algorithm 2) to progressively steer the downlink beam toward the active sensor based on the signal it has retro-directed. More precisely, at the $(k-1)$-th iteration, with $k\geq 1$, the AP transmits the interrogation signal with power $P_{\text{T}}$ using the current beamforming vector ${\bf{x}}[k-1]$ (step 2).
@@
This signal, after passing through the MIMO channel represented by the matrix ${\bf{H}}\in\mathbb{C}^{M\times N}$, is then received by the sensor (step 3), along with the AWGN noise ${\bm{\eta}}[k]\in\mathbb{C}^{1\times M}$. Thanks to the modulating SCM, the sensor reflects the received signal along the direction(s) of arrival, also changing its phase based on the data intended for the AP (step 4). Specifically, at the sensor side, information data is associated with the phase sequence $\left\{\theta[k]\right\}_{k=1}^{K}$, forming a packet of length $K$ symbols, according to any phase-based signaling scheme (e.g., BPSK). When reflecting the signal, the modulating SCM may also introduce a gain $g$, with $g<1$ if it is passive. In step 5, the AP receives the response ${\bf{y}}[k]\in\mathbb{C}^{1\times N}$ from the sensor, which incorporates the total AWGN at the receiver, ${\bf{n}}[k]\in\mathbb{C}^{1\times N}$, and accounts for the round-trip path experienced by the signal and the conjugation performed by the SCM through the matrix ${\bf{A}}=\sqrt{P_{\text{T}}}\,g\,{\bf{H}}^{{\mathrm{H}}}{\bf{H}}\in\mathbb{C}^{N\times N}$. Then, a normalized and conjugated version of the received vector ${\bf{y}}[k]$ is computed (step 6) and used as the updated beamforming vector ${\bf{x}}[k]$ in the subsequent iteration (see Fig. 13, second beam pattern).
Iteration after iteration, the beamforming vector is progressively refined over the air, ideally converging toward the optimal configuration (see Fig. 13, third beam pattern). Information data is also extracted by the AP at each iteration by correlating the current received vector with the previous beamforming vector ${\bf{x}}^{{\mathrm{H}}}[k-1]$, thus forming the scalar decision variable $u[k]$ (step 7). The decision on the modulation symbol conveyed by $\widehat{\theta}[k]$ at the $k$-th time instant is obtained by means of the function $\mathsf{detection(\cdot)}$, according to the adopted modulation scheme (step 8). It is worth noting that data demodulation occurs while the AP transmits the interrogation signal using the current beamforming vector, thanks to the full-duplex capability of the AP.
@@
In the absence of noise and data, the processing operated in Algorithm 2 corresponds to the well-known Power Method, which allows the estimate of the strongest eigenvector (top eigenvector) of a square matrix ${\bf{A}}$, described by the recurrence relation
${\bf{x}}[k]=\frac{{\bf{A}}\,{\bf{x}}[k-1]}{\left\|{\bf{A}}\,{\bf{x}}[k-1]\right\|}$ (45)
@@
being ${\bf{x}}[0]$ typically an initial random guess of the top eigenvector. For $k\rightarrow\infty$, the direction of ${\bf{x}}[k]$ converges to that of the top eigenvector. This means that ${\bf{x}}[k]$ tends to the direction of the top left eigenvector of the MIMO channel matrix ${\bf{H}}$, i.e., the optimal beamforming vector. In the presence of noise, new noise samples enter the loop at each iteration, and the convergence to the top eigenvector of the channel is no longer guaranteed. In this regard, the analysis in [17] shows that the iterative scheme still converges to the top eigenvector if the bootstrap SNR, defined as ${\mathsf{SNR}^{(\mathrm{boot})}=\mathsf{SNR}^{(\mathrm{max})}/N}$, is greater than one, where $\mathsf{SNR}^{(\mathrm{max})}$ denotes the maximum SNR achievable along the direction of the top eigenvector, which occurs when the AP beam is perfectly aligned with the sensor. When the algorithm operates successfully, the SNR experienced by the AP at the $k$-th iteration, denoted $\mathsf{SNR}[k]$, converges iteratively to $\mathsf{SNR}^{(\mathrm{max})}$ independently of the initial guess ${\bf{x}}[0]$.
@@
In Fig. 13, we also report some simulation results to investigate the performance of the scheme proposed in [17]. The scenario considered consists of an AP operating at 28 GHz and equipped with a uniform planar array of $N=20\times 20$ antennas deployed along the $xy$-plane and a sensor with a modulating SCM comprising $M=10\times 10$ cells assumed to lie on the same plane. Specifically, Fig. 13 shows the time evolution of $\mathsf{SNR}[k]$ under different conditions: The blue curve corresponds to the case ${\mathsf{SNR}^{(\mathrm{max})}}=35$ dB, associated with a bootstrap SNR of 5.45 dB, while the red curve corresponds to the case ${\mathsf{SNR}^{(\mathrm{max})}}=25$ dB, resulting in a bootstrap SNR of -5.45 dB. As shown, when ${\mathsf{SNR}^{(\mathrm{boot})}}$ exceeds 0 dB, $\mathsf{SNR}[k]$ converges to $\mathsf{SNR}^{(\mathrm{max})}$ after a few iterations, indicating that perfect alignment has been achieved and the packet can be successfully detected. Conversely, when ${\mathsf{SNR}^{(\mathrm{boot})}}$ is below 0 dB, $\mathsf{SNR}[k]$ converges to a much lower value and the link cannot be established.
It is worth noticing that the convergence time is below 10 time intervals, as confirmed by a more extensive investigation available in [17]. The algorithm is able to dynamically adjust the beam orientation in response to any change in the position of the sensor during ongoing communication, thereby exhibiting tracking capabilities. Notably, this is achieved without requiring analog-to-digital converter (ADC) chains at the UE, explicit channel estimation, and time-consuming beamforming or alignment processes. Application of this solution to grant-free random access schemes can be found in [58].

### v3

@@ V-C1 Reconfigurable Intelligent Surfaces
Specifically, by applying specific phase shifts to the incident EM waves, RISs can reflect signals in directions that deviate from the conventional Snell’s law. This phenomenon, referred to as the universal Snell’s law, allows RISs to impose arbitrary directionality on reflected signals, resulting in highly customizable beamforming and beam focusing. This capability enhances signal strength and reception, and depending on the specific hardware structure and reconfigurable elements of the RIS, additional functionalities such as beam collimation, beam splitting, polarization control, and various forms of analog processing can also be achieved [33, 32, 34]. This versatility positions RISs as a promising solution for addressing the complex challenges faced in modern wireless communication systems. Applications of RISs span communication, sensing, localization/tracking, and the detection of passive objects. Ongoing research is also investigating new avenues for leveraging RISs in mapping and imaging to reconstruct high-fidelity and high-resolution maps of the surrounding radio environment [54].
@@
As an example of the basic functionality of a RIS, let us consider a transmitting antenna, e.g., an access point (AP), that wants to transmit to a user equipment (UE) that is not in direct visibility due to obstruction raised by the presence of an obstacle, e.g., in a configuration similar to Fig. 11. Thanks to the deployment of the RIS, it is possible to create a virtual line-of-sight (LOS) link between the AP and the UE. Specifically, we consider a square RIS consisting of $K$ small unit cells (i.e., scattering elements) spaced $\lambda/2$ apart providing reconfigurable phase shifts, each being characterized by reflection coefficients $\rho_{i,j}=e^{\jmath\theta_{i,j}}$, with $i,j=1,2,\dots,\sqrt{K}$ being the 2D position indexes. According to the discussion in Section III-C, a conventional RIS with uncoupled elements is characterized by a diagonal reflection matrix ${\mathbf{R}}(\bm{\theta})={\rm diag}\left(r_{1},r_{2},\ldots,r_{K}\right)$ (diagonal RIS), where $r_{k}=\rho_{i,j}$ is the reflection coefficient of the $k$-th element at position $(i,j)$, where $k=i+\sqrt{K}(j-1)$, with $i,j=1,2,\dots,\sqrt{K}$.
@@
$\theta_{i,j}=-\pi i\left(u_{x}\left(\Theta_{\mathrm{inc}}\right)+u_{x}(\Theta_{\mathrm{ref}})\right)-\pi j\left(u_{y}\left(\Theta_{\mathrm{inc}}\right)+u_{y}(\Theta_{\mathrm{ref}})\right)\,$, (47)
@@
One important issue, often overlooked in the literature, arises from reflections in unwanted directions caused by uncontrolled RF sources in the environment. These reflections can generate interference, which may be challenging to mitigate. Diagonal RISs may be ineffective in addressing this problem due to the limited available number of DoF in their reflection matrix, which is equal to $K$. A potential alternative is the use of non-diagonal RISs, which allow controllable coupling between elements. This configuration significantly increases the available number of DoF, reaching up to K(K1)/2K\,(K-1)/2 in reciprocal structures [44]. An even greater number of DoF (up to $K^{2}$) can be achieved with non-reciprocal designs. Interestingly, a non-diagonal RIS, as a 2D structure, could theoretically achieve the same performance as a SIM, which is a 3D structure but with 2D input and output surfaces. However, as both technologies are still in their early stages, a detailed comparison in terms of implementation complexity is not yet available.
@@ V-C2 Self-Conjugating Metasurfaces
We now present an example of iterative over-the-air EM processing involving a transmitting/receiving MIMO AP and a scattering device located in their respective radiative regions. The AP interacts with the scattering device to retrieve information data embedded in the backscattered signal. From a SP perspective, the challenge is to achieve joint communication and optimal beamforming without requiring channel state information (CSI) estimation or signaling, thereby reducing the latency and overhead typically associated with large MIMO systems. The idea is to design the EM scattering device to perform the complex conjugation of the signal coming from the AP, thereby achieving retrodirectivity. Due to this processing capability, such a device is referred to as self-conjugating metasurface (SCM). Additionally, it is envisioned that the SCM introduces appropriate phase shifts, uniformly across all cells, into the retransmitted signals, thereby incorporating information through phase modulation.
This dual-function device, referred to as a modulating SCM in the following, can be considered an evolution of retrodirective antenna arrays, which are systems capable of reflecting incoming signals back toward their direction of arrival [55]. While retrodirective antennas have been studied for decades, the feasibility of retrodirective metasurfaces has only recently been demonstrated in the literature [56]. By leveraging modulating SCMs, it is possible to devise a communication scheme that involves an iterative exchange of signals between a full-duplex AP equipped with a conventional antenna array with $N$ elements, and a device equipped with a modulating SCM with $M$ elements, such as a sensor, which has queued data intended for the AP. The former continuously transmits an interrogation signal while the latter retro-directs the received signal after modulating it according to the data addressed to the interrogation device.
@@
0: Initialization: generate a guess unit norm beamforming vector ${\mathbf{x}}[0]$ ;
@@
   2: transmit: $\sqrt{P_{\text{T}}}\,{\mathbf{x}}[k-1]$ ;
@@
   3: ${\mathbf{z}}[k]=\sqrt{P_{\text{T}}}\,{\mathbf{H}}\,{\mathbf{x}}[k-1]+{\bm{\eta}}[k]$ ;
@@
   4: ${\mathbf{r}}[k]=ge^{\jmath\theta[k]}{\mathbf{z}}^{*}[k]$ ;
@@
   5: receive: ${\mathbf{y}}[k]=e^{\jmath\theta[k]}\,{\mathbf{A}}^{*}\,{\mathbf{x}}^{*}[k-1]+{\mathbf{n}}^{*}[k]$ ;
@@
   6: ${\mathbf{x}}[k]={{\mathbf{y}}^{*}[k]}/{\left\|{\mathbf{y}}[k]\right\|}$ ;
@@
The process begins (step 0 of the pseudo-code) with the AP generating a beamforming vector, ${\mathbf{x}}[0]\in\mathbb{C}^{1\times N}$. Since the position of the sensor is unknown at startup, this initial beamforming vector is chosen randomly (with norm one) to allow the signal transmitted using ${\mathbf{x}}[0]$ to cover multiple directions, as shown in Fig. 13, first beam pattern. An iterative process is then initiated (see Algorithm 2) to progressively steer the downlink beam toward the active sensor based on the signal it has retro-directed. More precisely, at the (k1)(k-1)-th iteration, with $k\geq 1$, the AP transmits the interrogation signal with power $P_{\text{T}}$ using the current beamforming vector ${\mathbf{x}}[k-1]$ (step 2).
@@
This signal, after passing through the MIMO channel represented by the matrix ${\mathbf{H}}\in\mathbb{C}^{M\times N}$, is then received by the sensor (step 3), along with the AWGN noise ${\bm{\eta}}[k]\in\mathbb{C}^{1\times M}$. Thanks to the modulating SCM, the sensor reflects the received signal along the direction(s) of arrival, also changing its phase based on the data intended for the AP (step 4). Specifically, at the sensor side, information data is associated with the phase sequence $\left\{\theta[k]\right\}_{k=1}^{K}$, forming a packet of length $K$ symbols, according to any phase-based signaling scheme (e.g., BPSK). When reflecting the signal, the modulating SCM may also introduce a gain $g$, with $g<1$ if it is passive. In step 5, the AP receives the response ${\mathbf{y}}[k]\in\mathbb{C}^{1\times N}$ from the sensor, which incorporates the total AWGN at the receiver, ${\mathbf{n}}[k]\in\mathbb{C}^{1\times N}$, and accounts for the round-trip path experienced by the signal and the conjugation performed by the SCM through the matrix ${\mathbf{A}}=\sqrt{P_{\text{T}}}\,g\,{\mathbf{H}}^{{\mathrm{H}}}{\mathbf{H}}\in\mathbb{C}^{N\times N}$. Then, a normalized and conjugated version of the received vector ${\mathbf{y}}[k]$ is computed (step 6) and used as the updated beamforming vector ${\mathbf{x}}[k]$ in the subsequent iteration (see Fig. 13, second beam pattern). Iteration after iteration, the beamforming vector is progressively refined over the air, ideally converging toward the optimal configuration (see Fig. 13, third beam pattern). Information data is also extracted by the AP at each iteration by correlating the current received vector with the previous beamforming vector ${\mathbf{x}}^{{\mathrm{H}}}[k-1]$, thus forming the scalar decision variable $u[k]$ (step 7). The decision on the modulation symbol conveyed by $\widehat{\theta}[k]$ at the $k$-th time instant is obtained by means of the function $\mathsf{detection(\cdot)}$, according to the adopted modulation scheme (step 8). It is worth noting that data demodulation occurs while the AP transmits the interrogation signal using the current beamforming vector, thanks to the full-duplex capability of the AP.
@@
In the absence of noise and data, the processing operated in Algorithm 2 corresponds to the well-known Power Method, which allows the estimate of the strongest eigenvector (top eigenvector) of a square matrix $\mathbf{A}$}, described by the recurrence relation
${\mathbf{x}}[k]=\frac{{\mathbf{A}}\,{\mathbf{x}}[k-1]}{\left\|{\mathbf{A}}\,{\mathbf{x}}[k-1]\right\|}\,$, (48)
@@
being ${\mathbf{x}}[0]$ typically an initial random guess of the top eigenvector. For $k\rightarrow\infty$, the direction of ${\mathbf{x}}[k]$ converges to that of the top eigenvector. This means that ${\mathbf{x}}[k]$ tends to the direction of the top left eigenvector of the MIMO channel matrix $\mathbf{H}$}, i.e., the optimal beamforming vector. In the presence of noise, new noise samples enter the loop at each iteration, and the convergence to the top eigenvector of the channel is no longer guaranteed. In this regard, the analysis in [17] shows that the iterative scheme still converges to the top eigenvector if the bootstrap SNR, defined as $\mathsf{SNR}^{(\mathrm{boot})}=\mathsf{SNR}^{(\mathrm{max})}/N$}, is greater than one, where $\mathsf{SNR}^{(\mathrm{max})}$ denotes the maximum SNR achievable along the direction of the top eigenvector, which occurs when the AP beam is perfectly aligned with the sensor. When the algorithm operates successfully, the SNR experienced by the AP at the $k$-th iteration, denoted $\mathsf{SNR}[k]$, converges iteratively to $\mathsf{SNR}^{(\mathrm{max})}$ independently of the initial guess ${\mathbf{x}}[0]$.
@@
In Fig. 13, we also report some simulation results to investigate the performance of the scheme proposed in [17]. The scenario considered consists of an AP operating at 28 GHz and equipped with a uniform planar array of $N=20\times 20$ antennas deployed along the $xy$-plane and a sensor with a modulating SCM comprising $M=10\times 10$ cells assumed to lie on the same plane. Specifically, Fig. 13 shows the time evolution of $\mathsf{SNR}[k]$ under different conditions: The blue curve corresponds to the case ${\mathsf{SNR}^{(\mathrm{max})}}=35\,\mathrm{dB}$, associated with a bootstrap SNR of $5.45\,\mathrm{dB}$, while the red curve corresponds to the case ${\mathsf{SNR}^{(\mathrm{max})}}=25\,\mathrm{dB}$, resulting in a bootstrap SNR of -5.45 dB. As shown, when $\mathsf{SNR}^{(\mathrm{boot})}$} exceeds $0\,\mathrm{dB}$, $\mathsf{SNR}[k]$ converges to $\mathsf{SNR}^{(\mathrm{max})}$ after a few iterations, indicating that perfect alignment has been achieved and the packet can be successfully detected. Conversely, when $\mathsf{SNR}^{(\mathrm{boot})}$} is below $0\,\mathrm{dB}$, $\mathsf{SNR}[k]$ converges to a much lower value and the link cannot be established.
It is worth noticing that the convergence time is below 10 time intervals, as confirmed by a more extensive investigation available in [17]. The algorithm is able to dynamically adjust the beam orientation in response to any change in the position of the sensor during ongoing communication, thereby exhibiting tracking capabilities. Notably, this is achieved without requiring ADC chains at the sensor, explicit channel estimation, and time-consuming beamforming or alignment processes. Application of this solution to grant-free random access schemes can be found in [57].